OpenStudy (anonymous):

You are given that f(x) = x^3 - 19x + 30 1.) Calculate f(0) and f(3). Hence write down a factor of f(x). Thanks.

OpenStudy (anonymous):

@amistre64 @Hero

OpenStudy (anonymous):

can u evaluate f(0) and f(3) ?

OpenStudy (anonymous):

What does evaluate mean?

OpenStudy (anonymous):

calculate-compute-

OpenStudy (anonymous):

Hmmm hang on. Do you mean put them into the f(X)= thing?

hero (hero):

Evaluate means to do this: f(x) = x^3 - 19x + 30 f(0) = (0)^3 - 19(0) + 30 = ? f(3) = (3)^3 - 19(3) + 30 = ?

OpenStudy (anonymous):

Yes ok Ill have a go at that and tell you what I got thanks.

OpenStudy (anonymous):

Right for f(0) I got 30 and the next one 0

hero (hero):

f(3) = 0 Are you sure?

OpenStudy (anonymous):

Ill do it again

OpenStudy (anonymous):

Did it again and got 0 again.

hero (hero):

f(3) = (3)^3 - 19(3) + 30 = 27 - 57 + 30 = 30 + 27 - 57 = 57 - 57 = 0 OK

OpenStudy (anonymous):

Yes!! So now I have two more questions. Are you up for helping?

OpenStudy (anonymous):

Wait what is the 'hence write down the factor of f(x)' bit?

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@estudier

hero (hero):

If you calculated f(3) = 0, then (x-3) is a factor

OpenStudy (anonymous):

:) I remeber that now. The 0 ones it is.

OpenStudy (anonymous):

(ii) Find p and q such that f(x) = (x-2)(x^2 + px +q). How does one go about this?

OpenStudy (anonymous):

@karatechopper @myininaya

OpenStudy (anonymous):

@Coolsector

OpenStudy (anonymous):

(x-2)(x^2 + px +q) = x^3 +px^2 +qx -2x^2 -2px -2q = x^3 +(p-2)x^2 +(q-2p)x - 2q and the orig is x^3 - 19x + 30 so we know that p-2 should be 0 and q-2p should be -19 and -2q should be 30 can you answer now ?

OpenStudy (anonymous):

Wait....Where has that big line came from? ohhh you timesed out the brackets! Ok wait a min.

OpenStudy (anonymous):

I get it up to when you start saying the so we know p-2=0 bit. I dont know how we know that.

OpenStudy (anonymous):

ok you see we have a term (p-2)x^2 but in the general equation we dont have x^2 at all .. so we have to make it disappear .. we do this by saying p-2 = 0